OpenStudy (anonymous):

Which exponential function goes through the points (1, 16) and (4, 128)? (4 points)

OpenStudy (anonymous):

please help

OpenStudy (dmndlife24):

What are the answer choices?

OpenStudy (anonymous):

it doent have any

OpenStudy (dmndlife24):

it's asking for "which exponential function"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (dmndlife24):

found this: http://inquus.org/updates/52d764b7e4b0e63d49ac645b

OpenStudy (dmndlife24):

So it must be f(x) = 8(2)x

OpenStudy (anonymous):

thanks

OpenStudy (dmndlife24):

Your welcome:)